# Evolution of the pseudogap band structure in a system of electron-correlated lattice polarons

### v1

@@ I Introduction
The addition of strong electron-phonon couplings to a system with strong electron correlations results in a modulation of the competition between the kinetic energy of the carriers and their Coulomb interaction energy, all set against the backdrop of lattice polaron effects. The intricate interplay of interactions gives rise to a region in the phase diagram where a unique type of quasiparticle exists – electron-correlated lattice polarons. The properties and conditions for their formation differ significantly from those of polarons in uncorrelated systems and are, among other things, influenced by the competition between various types of electron-phonon coupling mechanisms [34, 35]. The first one is associated with changes in the local energy of charge carriers and is described by the Holstein Hamiltonian. The second corresponds to the modulation of the hopping parameters and is usually referred to as the Peierls contribution or the transitive electron-phonon interaction. A proper description of the phase diagram of a polaron system requires taking both types of microscopic mechanisms of electron-phonon coupling into account. This conclusion is supported by previous studies addressing a range of issues, including the non-analytic behavior of polaron characteristics [36, 37, 38, 39, 30], crossovers between polaron and bipolaron transformations [34], nontrivial topological effects [40, 41], and antiferromagnetic order induced by transitive electron-phonon interaction [42, 43, 44].
@@ II Model and calculation method
Here, the vector $\bf{g}$ enumerates the sites of a square lattice centered on the atoms with partially filled $d$-orbitals, $\alpha$ is the orbital index, and $\alpha=p,d$, with $p$ corresponding to the ligand atoms, ${\bf{g}}_{\alpha}={\bf{g}}_{l}$ for $\alpha=p$, where vector ${\bf{g}}_{l}$ runs through the positions of the ligand atoms in the unit cell. The values $\varepsilon_{\alpha}$ determine the local atomic energies with respect to the chemical potential $\mu$. The operators $d_{\bf{g},\sigma}^{\dagger}$ and $p_{{\bf{g}}_{l},\sigma}^{\dagger}$ create a hole with spin $\sigma$ in the $d$- or $p$-orbitals, respectively, ${n_{{\bf{g}},\sigma}^{\alpha}}$ are the particle number operators, $f_{\bf{g}}^{\dagger}$ and $f_{\bf{g}}$ are the operators of creation and annihilation of the breathing planar mode with frequency ${\omega_{ph}}$. The parameters ${U_{\alpha}}$ and ${V_{pd}}$ are the matrix elements of Coulomb repulsion, ${t_{pd}}$ and ${t_{pp}}$ are the overlap integrals of corresponding orbitals, $-\sigma=\bar{\sigma}$, $P_{pp}$ and $P_{pd}$ are the phase factors equal to either $1$ or $-1$, depending on whether the orbitals with real wave functions have the same or opposite signs in the overlap region.
@@
The results presented below have been obtained within the polaronic version [30, 45] of the generalized tight-binding method (pGTB), which is essentially a variant of the cluster perturbation theory. A distinctive feature of the approach is the introduction of an adequate set of quasiparticle excitations that are formed by simultaneously accounting for both short-range electron-electron correlations and electron-lattice coupling within an effective cluster. This makes it possible to model solutions across a fairly wide range of electron-lattice coupling parameters in a correlated system. Following the ideology of the GTB method, we carry out canonical transformations of the Hamiltonian, which allows us to separate equation (II) into intra-cluster and inter-cluster contributions. Through exact diagonalization of the cluster Hamiltonian, we obtain its many-body eigenstates. The interactions between clusters that lead to the formation of dispersion relations for quasiparticle excitations are then analyzed using perturbation theory. To consider the inter-cluster contributions, we employ projection techniques in the equations of motion for thermodynamic two-time Green’s functions [47, 48, 34] within a generalized Hartree-Fock approximation that takes into account the interaction between charge carriers and spin fluctuations.
@@ III.1 The competition of Coulomb and Holstein interactions
To determine the second critical point along the $\lambda_{H}$ axis, we compare the on-site energies of the polaron $-\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$ and the bipolaron $U_{d}-2\Delta-2\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$, shifting the energy scale for convenience so that the condition $\varepsilon_{d}=0$ is satisfied. It is evident that in the limit of strong electron correlations, Coulomb repulsion prevents the formation of bipolarons, thereby preserving the dominance of polaron formations. This situation changes with an increase in the strength of electron-phonon coupling. The formation of a local bipolaronic state becomes more favorable since only $-\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}\geq U_{d}-2\Delta-2\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$, from which we immediately obtain that $\lambda^{\left({2}\right)}_{H}=U_{d}/W\left(2+\langle{n_{{\bf{g},\sigma}}^{d}}\rangle\right)$.
@@ III.2 The essential role of Peierls’ contribution
Below, the role of lattice polarons in the formation of a pseudogap will be examined in the range of electron-phonon interaction parameters that describe the system prior to bipolaron transformations. For model (II), this is the region of the phase diagram in Fig. 1, located below the red curve. We will show how qualitative changes in the electronic structure evolve at the Fermi level with the varying strength of the electron-lattice interaction and consider their evolution with doping.
@@ IV Analysis of electronic spectra
In the limit of strong and intermediate electron correlations, the pseudogap state of a system without electron-phonon interactions emerges against the background of fluctuations in short-range antiferromagnetic order. This leads to an anisotropic rearrangement of the electron band strusture, which evolves with doping from a non-Fermi liquid in lightly doped systems to a typical Fermi liquid in overdoped systems [2, 52, 53, 54, 55], Fig. 2 (a-g). The transition between different regimes occurs through two quantum critical points [54], caused by changes in the Fermi surface topology at doping level $x_{c1}\approx 0.18$ and $x_{c2}\approx 0.27$. The first critical point corresponds to the merging of small hole pockets formed near the points (±$\left(\pm{\pi}/{2};\pm{\pi}/{2}\right)$ and is accompanied by a logarithmic Van Hove singularity in the density of states. The second critical point corresponds to the collapse of the inner contour of the Fermi surface around the point $\left({\pi},{\pi}\right)$ and is accompanied by a Heaviside step-type singularity in the density of states.
@@
The formation of a narrow, partially flat polaron band begins around the point $\left({\pi},{\pi}\right)$ and is accompanied by the emergence of an additional peak with lower intensity in the antinodal direction of the spectral function. In order to understand how the parameters of electron-lattice coupling and doping control the effects of the narrow polaron band, we performed a detailed study of the phase diagram of the system: the minimum step sizes for the values of $\lambda_{H}$ and $\lambda_{P}$ were $0.005$, while the minimum interval for doping was $\Delta x=0.01$. It was found that (i) the narrow polaron band aligns with the Fermi level in the region of the EPI parameters describing the correlated polaron, or approaching this region (that is, near and above the red line corresponding to the polaron, but below the black line indicating the bipolaron transformations, to the left of the point of their intersection in Fig. 1); (ii) this occurs in the non-Fermi-liquid regime at concentrations of doped holes that are less than or approximately equal to the first critical value $x_{c1}$.
@@
Qualitative changes in the Fermi surface resulting from polaron effects are clarified in Figure 11. The insets (a) and (c) present a comparison between the band structure of a non-modulated system and a crucial stage of a strongly modulated one. In the system characterized by strong EPI, the partially flat polaron band, formed around the point $\left({\pi},{\pi}\right)$, overlaps with the original top of the valence band lying below it at the points (±$\left({\pm\pi}/{2};{\pm\pi}/{2}\right)$. This overlapping further extends to the original band structure near the $X$ and $Y$. If the Fermi level lies within the energy range from the top of the newly formed hybrid band to the level corresponding to the top of the untransformed band, an arc-like Fermi surface is observed. In the corresponding part of the spectrum, the density of states $N\left(\omega\right)$ exhibits a feature in the form of a shoulder, which leads to the maximum of the function $N_{F}\left(x\right)$.

### v2

@@ I Introduction
The addition of strong electron-phonon couplings to a system with strong electron correlations results in a modulation of the competition between the kinetic energy of the carriers and their Coulomb interaction energy, all set against the backdrop of lattice polaron effects. The intricate interplay of interactions gives rise to a region in the phase diagram where a unique type of quasiparticle exists – electron-correlated lattice polarons. The properties and conditions for their formation differ significantly from those of polarons in uncorrelated systems and are, among other things, influenced by the competition between various types of electron-phonon coupling mechanisms [34, 35]. The first one is associated with changes in the local energy of charge carriers and is described by the Holstein Hamiltonian. The second corresponds to the modulation of the hopping parameters and is usually referred to as the Peierls contribution or the transitive electron-phonon interaction. Notably, the latter type of electron-lattice coupling has drawn significant attention since the seminal papers of Baris̆ić, Labbé, and Friedel on transition-metal superconductivity [36, 37, 38] and later by Su, Schrieffer, and Heeger on soliton formation in conducting polymers [39]. A proper description of the phase diagram of a polaron system requires taking both types of microscopic mechanisms of electron-phonon coupling into account. This conclusion is supported by previous studies addressing a range of issues, including the non-analytic behavior of polaron characteristics [40, 41, 42, 43, 30], crossovers between polaron and bipolaron transformations [34], nontrivial topological effects [39, 44], and antiferromagnetic order induced by transitive electron-phonon interaction [45, 46, 47].
@@ II Model and calculation method
Here, the vector $\bf{g}$ enumerates the sites of a square lattice centered on the atoms with partially filled $d$-orbitals, $\alpha$ is the orbital index, and $\alpha=p,d$, with $p$ corresponding to the ligand atoms, ${\bf{g}}_{\alpha}={\bf{g}}_{l}$ for $\alpha=p$, where vector ${\bf{g}}_{l}$ runs through the positions of the ligand atoms in the unit cell. The values $\varepsilon_{\alpha}$ determine the local atomic energies with respect to the chemical potential $\mu$. The operators $d_{\bf{g},\sigma}^{\dagger}$ and $p_{{\bf{g}}_{l},\sigma}^{\dagger}$ create a hole with spin $\sigma$ in the $d$- or $p$-orbitals, respectively, $n_{{\bf{g}},\sigma}^{\alpha}$} are the particle number operators, $f_{\bf{g}}^{\dagger}$ and $f_{\bf{g}}$ are the operators of creation and annihilation of the breathing planar mode with frequency $\omega_{ph}$}. The parameters $U_{\alpha}$} and $V_{pd}$} are the matrix elements of Coulomb repulsion, $t_{pd}$} and $t_{pp}$} are the overlap integrals of corresponding orbitals, $\sigma=\bar{\sigma}$, $P_{pp}$ and $P_{pd}$ are the phase factors equal to either $1$ or 1-1, depending on whether the orbitals with real wave functions have the same or opposite signs in the overlap region.
@@
The results presented below have been obtained within the polaronic version [30, 48] of the generalized tight-binding method (pGTB) [50], which is essentially a variant of the cluster perturbation theory. A distinctive feature of the approach is the introduction of an adequate set of quasiparticle excitations that are formed by simultaneously accounting for both short-range electron-electron correlations and electron-lattice coupling within an effective cluster. This makes it possible to model solutions across a fairly wide range of electron-lattice coupling parameters in a correlated system. Following the ideology of the GTB method, we carry out canonical transformations of the Hamiltonian, which allows us to separate equation (II) into intra-cluster and inter-cluster contributions. Through exact diagonalization of the cluster Hamiltonian, we obtain its many-body eigenstates. The interactions between clusters that lead to the formation of dispersion relations for quasiparticle excitations are then analyzed using perturbation theory. To consider the inter-cluster contributions, we employ projection techniques in the equations of motion for thermodynamic two-time Green’s functions [51, 52, 34] within a generalized Hartree-Fock approximation that takes into account the interaction between charge carriers and spin fluctuations.
@@ III.1 The competition of Coulomb and Holstein interactions
To determine the second critical point along the $\lambda_{H}$ axis, we compare the on-site energies of the polaron $\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$ and the bipolaron $U_{d}-2\Delta-2\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$, shifting the energy scale for convenience so that the condition $\varepsilon_{d}=0$ is satisfied. It is evident that in the limit of strong electron correlations, Coulomb repulsion prevents the formation of bipolarons, thereby preserving the dominance of polaron formations. This situation changes with an increase in the strength of electron-phonon coupling. The formation of a local bipolaronic state becomes more favorable since only $\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}\geq U_{d}-2\Delta-2\Delta{\langle{n_{{\bf{g},\sigma}}^{d}}\rangle}$, from which we immediately obtain that $\lambda^{\left({2}\right)}_{H}=U_{d}/W\left(2+\langle{n_{{\bf{g},\sigma}}^{d}}\rangle\right)$.
@@ III.2 The essential role of Peierls’ contribution
Below, the role of lattice polarons in the formation of a pseudogap will be examined in the range of electron-phonon interaction parameters that describe the system prior to bipolaron transformations. For model (II), this is the region of the phase diagram in Fig. 1, located below the red curve. Within the pGTB approach, we will demonstrate how qualitative changes in the electronic structure evolve at the Fermi level with the varying strength of the electron-lattice interaction and consider their evolution with doping.
@@ IV Analysis of electronic spectra
In the limit of strong and intermediate electron correlations, the pseudogap state of a system without electron-phonon interactions emerges against the background of fluctuations in short-range antiferromagnetic order. This leads to an anisotropic rearrangement of the electron band strusture, which evolves with doping from a non-Fermi liquid in lightly doped systems to a typical Fermi liquid in overdoped systems [2, 56, 57, 58, 59], Fig. 2 (a-g). The transition between different regimes occurs through two quantum critical points [58], caused by changes in the Fermi surface topology at doping level $x_{c1}\approx 0.18$ and $x_{c2}\approx 0.27$. The first critical point corresponds to the merging of small hole pockets formed near the points $\left(\pm{\pi}/{2};\pm{\pi}/{2}\right)$ and is accompanied by a logarithmic Van Hove singularity in the density of states. The second critical point corresponds to the collapse of the inner contour of the Fermi surface around the point $\left({\pi},{\pi}\right)$ and is accompanied by a Heaviside step-type singularity in the density of states.
@@
The formation of a narrow, partially flat polaron band begins around the point $\left({\pi},{\pi}\right)$ and is accompanied by the emergence of an additional peak with lower intensity in the antinodal direction of the spectral function. In order to understand how the parameters of electron-lattice coupling and doping control the effects of the narrow polaron band, we performed a detailed study of the phase diagram of the system (see Appendix A).
It was found that (i) the narrow polaron band aligns with the Fermi level in the region of the EPI parameters describing the correlated polaron, or approaching this region (that is, near and above the black line corresponding to the polaron, but below the red line indicating the bipolaron transformations, to the left of the point of their intersection in Fig. 1); (ii) this occurs in the non-Fermi-liquid regime at concentrations of doped holes that are less than or approximately equal to the first critical value $x_{c1}$.
@@
Qualitative changes in the Fermi surface resulting from polaron effects are clarified in Figure 11. The insets (a) and (c) present a comparison between the band structure of a non-modulated system and a crucial stage of a strongly modulated one. In the system characterized by strong EPI, the partially flat polaron band, formed around the point $\left({\pi},{\pi}\right)$, overlaps with the original top of the valence band lying below it at the points $\left({\pm\pi}/{2};{\pm\pi}/{2}\right)$. This overlapping further extends to the original band structure near the $X$ and $Y$. If the Fermi level lies within the energy range from the top of the newly formed hybrid band to the level corresponding to the top of the untransformed band, an arc-like Fermi surface is observed. In the corresponding part of the spectrum, the density of states $N\left(\omega\right)$ exhibits a feature in the form of a shoulder, which leads to the maximum of the function $N_{F}\left(x\right)$.
@@ Acknowledgements.
Appendix A Principles and Practical Implementation of the pGTB Method
The polaronic version of the generalized tight-binding method (p-GTB), as well as the original GTB, belongs to the cluster perturbation theory framework. The approach begins with the decomposition of the full Hamiltonian (II) into intracluster and intercluster components $H=H_{c}+H_{cc}$, where the intracluster term $H_{c}=\sum\nolimits_{\bf{g}}{H_{\bf{g}}}$ is summed over all lattice sites $\bf{g}$}. The problem is addressed via canonical transformations of the Hamiltonian, with the transformation matrices depending on the unit cell symmetry, the irreducible representations of its point group, and the symmetry properties of the phonon modes [61, 62, 63, 64, 48]. The constructed unit cell cluster captures the realistic structure and chemical bonding environment of the planar tetragonal system, while the incorporated full-breathing mode preserves its point symmetry.
In the new symmetric cell representation, the strong on-site localization of Wannier orbitals induces the renormalization of all parameters of the full Hamiltonian. For example, the parameters $U_{p}$ and $V_{pd}$ of the Coulomb interactions related to ligand holes transform as ${U_{p}}{\Psi_{gfmn}}$ and $V_{pd}{\Phi_{gfm}}$, respectively. The intracell coefficients are $\Psi_{0000}=0.21$ and $\Phi_{000}=0.918$, while the intercell terms are $\Psi_{0001}=-0.03$, $\Phi_{001}=0.13$, and $\Phi_{002}=-0.02$. Thus, the Coulomb interactions are strongly localized within the unit cell of $H_{\bf{g}}$, allowing the intercell contributions from $U_{p}$ and $V_{pd}$ to be neglected. Generally, this approach introduces a methodological constraint: the retained intercluster electron-phonon contributions must dominate over the discarded intercluster Coulomb interaction terms.
The renormalized electron-phonon coupling parameters $M_{{\bf{gg}}^{\prime}}^{d}={M_{d}}{\mu_{{\bf{gg}}^{\prime}}}$ and $M_{{\bf{g}\bf{g}^{\prime}\bf{h}}}^{bd}=-2{M_{pd}}{\mu_{{\bf{gg}}^{\prime}}}{\mu_{{\bf{hg}}^{\prime}}}$ also exhibit rapid distance decay, as quantified by $\mu_{00}=0.96$, $\mu_{01}=-0.14$, and $\mu_{11}=-0.02$. Despite their short-range nature, these intercluster contributions remain essential in certain physical contexts, for example, in determining the dispersion of phonon excitations [34]. Furthermore, these terms restore the full Hamiltonian’s translational symmetry when modified by a vibrational phase factor accounting for inter-site phase differences. To date, the simultaneous accurate treatment of both long-range and short-range correlations in systems with strong interactions remains an unsolved problem. We assume that in the strong-coupling regime, the key features of small-polaron electronic structure and the density of states are predominantly determined by the competition between localized electron-electron and electron-phonon interactions. Consequently, the aforementioned intercluster terms, including phase correlations, have been neglected in our analysis.
The complete expression for the Hamiltonian (II) in its intermediate representation, $H_{c}+H_{cc}$, is given in Ref. [34]. The next step involves the exact diagonalization of the effective Hamiltonian for the unit cell cluster described by the Hamiltonian $H_{\bf{g}}$. The procedure is performed for each system configuration characterized by $n_{h}$ holes per unit cell and followed by a controlled truncation of high-energy states with minimal spectral weight. Our truncation of the phonon Hilbert space ensures that the ground state and first excited states energies converge with less than $1\%$ error for all cases (0, 1, and 2 holes per site) while preserving all essential spectral features.
Using the complete set of multielectron-multiphonon eigenstates $\left|q\right\rangle$} obtained through exact diagonalization of $H_{\mathbf{g}}$, we express the full Hamiltonian in terms of Hubbard X-operators. Indeed, any operator $A_{\bf{g}}$ can be presented as a linear combination:
$A_{\bf{g}}=\sum\limits_{q,q^{\prime}}{\left\langle q\right|A_{\bf{g}}\left|{q^{\prime}}\right\rangle{X_{\bf{g}}^{q,q^{\prime}}}}=\sum\limits_{Q}{{\gamma}_{A,Q}{X_{\bf{g}}^{Q}}},$ (3)
where the index $q$ encompasses all relevant quantum numbers and ${\gamma_{A,Q}}=\left\langle{q\left|{{A_{\bf{g}}}}\right|q^{\prime}}\right\rangle$ is the matrix element for the transition from the initial state $\left|q^{\prime}\right\rangle$ to the final state $\left|q\right\rangle$, i.e. for a pair of states $Q=\left({q,q^{\prime}}\right)$, under the action of the operator $A$. The Hamiltonian then reduces to the compact form:
$H_{c}=\sum\limits_{{\bf{g}},q}{{E_{q}}{X_{\bf{g}}^{q,q}}},$ (4)
${H_{cc}}=\sum\limits_{{{\bf{g}}\neq{\bf{g}}^{\prime}\neq{\bf{g^{\prime\prime}}}}}{\sum\limits_{Q,Q^{\prime}}{\left({T_{{\bf{gg^{\prime}}}}^{Q^{\prime}Q}{\delta_{{\bf{g^{\prime\prime}g^{\prime}}}}}\overset{\dagger}{\mathop{X_{\mathbf{g}}^{Q^{\prime}}}}X_{{\bf{g^{\prime}}}}^{Q}+}\right.}}$
$+\sum\limits_{Q^{\prime\prime}}{\left({M_{{\bf{gg^{\prime}}}}^{Q^{\prime\prime}Q^{\prime}}{\delta_{{\bf{g^{\prime\prime}g^{\prime}}}}}{\delta_{Q^{\prime}Q}}+M_{{\bf{gg^{\prime}}}{\bf{g^{\prime\prime}}}}^{Q^{\prime\prime}Q^{\prime}Q}}\right)\left.{X_{\bf{g}}^{Q^{\prime\prime}}\overset{\dagger}{\mathop{X_{\mathbf{g^{\prime}}}^{Q^{\prime}}}}X_{\bf{g^{\prime\prime}}}^{Q}}\right)},$ (5)
with coefficients as
$T_{{\bf{gg^{\prime}}}}^{Q^{\prime}Q}={\tilde{t}}_{{\bf{gg^{\prime}}}}^{pp}\gamma_{{\tilde{p}}_{\sigma},Q^{\prime}}^{\dagger}{\gamma_{{{\tilde{p}}_{\sigma}},Q}}+{\tilde{t}}_{{\bf{gg^{\prime}}}}^{pd}\Gamma_{Q^{\prime}Q}^{pd},$
$M_{{\bf{gg^{\prime}}}}^{Q^{\prime\prime}Q^{\prime}}={\tilde{M}}_{{\bf{gg^{\prime}}}}^{d}\Gamma_{Q^{\prime\prime}}^{\tilde{f}}\gamma_{{d_{\sigma}},Q^{\prime\prime}}^{\dagger}{\gamma_{{d_{\sigma}},Q^{\prime}}},$
$M_{{\bf{gg^{\prime}}}{\bf{g^{\prime\prime}}}}^{Q^{\prime\prime}Q^{\prime}Q}={\tilde{M}}_{{\bf{gg^{\prime\prime}g^{\prime}}}}^{pd}\Gamma_{Q^{\prime\prime}}^{\tilde{f}}\Gamma_{Q^{\prime}Q}^{pd},$
$\Gamma_{Q^{\prime}Q}^{pd}=\gamma_{d_{\sigma},Q^{\prime}}^{\dagger}{\gamma_{{{\tilde{p}}_{\sigma}},Q}}+\gamma_{{\tilde{p}}_{\sigma},Q^{\prime}}^{\dagger}{\gamma_{{d_{\sigma}},Q}},$
$\Gamma_{Q^{\prime\prime}}^{\tilde{f}}=\gamma_{{\tilde{f}},Q^{\prime\prime}}^{\dagger}+{\gamma_{\tilde{f},Q^{\prime\prime}}}.$
[…]
It should be emphasized that the employed mathematical formalism accurately captures the essential physics of band structure formation in correlated systems, particularly the spectral weight redistribution among many-body quasiparticle excitations arising from the local competition between strong electron-electron and electron-lattice interactions. The intensity of these quasiparticle excitations is governed by the overlap of the matrix elements of the participating states and their occupation numbers. Consequently, the redistribution of spectral weights between quasiparticle bands depends on doping, temperature, and the relative strengths of electron-phonon coupling and Coulomb interactions.
Within the Hubbard operator approach, the retarded two-time single-particle Green’s function expands into a series of quasiparticle propagators
$G_{\mathbf{gg^{\prime}},\sigma}^{\alpha\alpha^{\prime}}\left(t,t^{\prime}\right)=$ (6)
[…]
where $a_{\alpha,\mathbf{g},\sigma}$ represents the annihilation operator for a hole with spin $\sigma$ at a lattice site $\mathbf{g}$ on the orbital $\alpha$ and the indices $Q_{h}$ enumerate all allowed transitions between pairs of eigenstates where the hole number decreases by one in the final state.
To determine the quasiparticle band structure of correlated holes coupled to phonons, we employ [34] the projection operator method in the equation of motion for the matrix Green’s function $\hat{G}_{\mathbf{gg^{\prime}},\sigma}^{{{Q}_{h}}{{Q}_{h^{\prime}}}}\left(t,t^{\prime}\right)$. In the generalized mean-field approximation, which incorporates the interaction of charge carriers with spin fluctuations, the quasiparticle spectrum is described by the matrix ${{{\hat{\omega}}}_{\mathbf{k},\sigma}}^{{{Q}_{h}}{{Q}_{h^{\prime}}}}$, with elements given by
$\omega_{\mathbf{k},\sigma}^{{{Q}_{h}}{{Q}_{h^{\prime}}}}=F_{H}^{{{Q}_{h}}}F_{H}^{{{Q}_{h^{\prime}}}}T_{\mathbf{k}}^{{{Q}_{h}}{{Q}_{h^{\prime}}}}\pm\sum\limits_{\mathbf{q}}{{{c}_{\mathbf{q}}}T_{\mathbf{k}-\mathbf{q}}^{{{Q}_{h}}{{Q}_{h^{\prime}}}}}$ (7)
$+F_{H}^{{{Q}_{h}}}{{E}_{{{Q}_{h}}}}{{\delta}_{{{Q}_{h}},{{Q}_{h^{\prime}}}}}.$
The plus (minus) sign is taken when the transitions $Q_{h}$ and $Q_{h^{\prime}}$ belong to the same (different) subspaces of the Hilbert space, ${\delta}_{{{Q}_{h}},{{Q}_{h^{\prime}}}}$ is the Kronecker delta-symbol, $F_{H}^{{{Q}_{h}}}=\left\langle{{X}_{\mathbf{g}}^{q,q}+{X}_{\mathbf{g}}^{q^{\prime},q^{\prime}}}\right\rangle$ for index ${Q}_{h}=\left({q,q^{\prime}}\right)$, and ${E}_{{Q}_{h}}=\left({E_{q^{\prime}}-E_{q}}\right)$. Fourier transform of the static spin correlation function is given by expression ${c_{\bf{q}}}=\sum\limits_{\left({{\bf{g}}-{\bf{g}}^{\prime}}\right)}{{c_{{\bf{gg}}^{\prime}}}\exp\left[{-{\bf{q}}\left({{\bf{g}}-{\bf{g}}^{\prime}}\right)}\right]}$, where ${c_{{\bf{gg}}^{\prime}}}=3\left\langle{S_{\bf{g}}^{z}S_{{\bf{g}}^{\prime}}^{z}}\right\rangle$. It characterizes [65] the doping-dependent evolution of spin-liquid properties of underdoped cuprates.
The present implementation of the method correctly describes short-range correlations of competing electron-electron and electron-phonon interactions. This enables analysis of small-polaron electronic structure evolution in the normal state. However, to describe phenomena such as superconducting states, large-to-small polaron transitions, or large polaron properties, the inclusion of relevant long-range correlations is required.
To systematically explore the phase diagram of the system, we calculate the electronic band structure using adaptive resolution for electron-phonon couplings ($\lambda_{H}$, $\lambda_{P}$) and variable doping intervals $\Delta x$. In parameter-sensitive regions where spectral properties exhibit substantial variations, we employ minimum step sizes of $0.05$ for $\lambda_{H}$ and $0.005$ for $\lambda_{P}$, with a doping concentration resolution of $0.01$. Specifically, for the system with $U_{d}=9$ (defined in Section II), we conducted a detailed investigation of the parameter space region $0.8\leq\lambda_{H}\leq 1.6$ and $0\leq\lambda_{P}\leq 0.08$. For each parameter combination, we recompute the complete electronic band structure, including the chemical potential. Our pGTB simulations adopted an $400\times 400$ k-point mesh for full Brillouin zone sampling. All calculations were performed on high-performance computing clusters using parallelized algorithms. For the parameter range considered in Section IV, the average execution time was $15$ minutes per phase diagram point. We note that the computational cost increases substantially with decreasing phonon frequency $\omega_{ph}$ or as the electron-phonon coupling $\lambda_{H}$ approaches its critical value $\lambda_{H}^{\left(3\right)}$. In the present work, we do not examine the band structure in these parameter regimes. Furthermore, our analysis does not include interactions with Jahn-Teller active vibrational modes that lift the orbital degeneracy of ground electronic states, nor temperature-dependent effects, which may become significant [48, 60] in polaron systems due to Frank-Condon resonance excitations.